OpenStudy (anonymous):

How can 2x2 = x2 + 4x + 8 be set up as a system of equations?

OpenStudy (anonymous):

y = 2x2 y = x2 + 4x + 8

OpenStudy (anonymous):

thank you (:

OpenStudy (anonymous):

It's C......... Out of your choices! NO Problem! If you have another question just mention me!

OpenStudy (anonymous):

can you help me on one more please <3 @cowgirlcrazy66

OpenStudy (anonymous):

Sure!

OpenStudy (anonymous):

A company sells boxes of cow bells (b) for $20 and boxes of air horns (h) for $30. They can make a batch of cow bells that fills 8 boxes or a batch of air horns that fills 9 boxes. The company only has 53 boxes. Which system of equations helps the company plan to make $275? b = 53 − h 8b + 9h = 53 8b = 53 − 9h 20b + 30h = 275 b = 53 − h 20b + 30h = 275 8b = 53 − 9h 20b − 30h = 275 @cowgirlcrazy66

hero (hero):

There really wasn't a way to explain the first question without giving the answer. @GloriaSCastillo, you should familiarize yourself more with solving systems of equations.

OpenStudy (anonymous):

@Hero yeah i know what you mean thank you though

hero (hero):

@GloriaSCastillo what are the given variables for this second problem?